# Dynamical stability of various convex graphical translators

### v1

@@ 1 Introduction
Ecker and Huisken [undefn], [undefm] showed by maximum principle arguments that when the initial hypersurface $M_{0}$ is an entire graph of a locally Lipchitz function, then there exists a longtime smooth solution to mean curvature flow starting from $M_{0}$ which its time slices are also entire graphs. Furthermore, they showed that if $M_{0}$ is ‘straight at infinity’, then after suitable rescaling, the hypersurfaces converge to a self expander, a special type of hypersurfaces which move by expansion under mean curvature flow. Later, Clutterbuck [undefi] (see also [undefj]) relaxed local Lipchitz continuity to just continuity. Uniqueness of entire graph solutions starting from locally Lipchitz, proper graph was proved by Daskalopolous and Sáez [undefl].
In the case when $\Omega\subset\mathbf{R}^{n}$ is a general domain, Sáez and Schnürer [undefaa] showed the existence of solutions to mean curvature flow when the initial data is a graph of a locally Lipchitz, proper function over $\Omega$. They further showed that such solutions remain graphical over evolving domain.
Our first result is the longtime existence of a graphical solution to mean curvature flow starting from a graph of a continuous function defined over $\Omega^{n}_{b}=\mathbf{R}^{n-1}\times(-b,b)$ for some $b>0$, analogous to results in [undefn], [undefi].
@@
Compared to results in [undefaa], although we are only working with $\mathbf{R}^{n-1}\times(-b,b)$ instead of general domains, our result relaxes the local Lipchitz continuity to just continuity. We are also assuming a slightly weaker condition on $u_{0}$ than it being a proper function in view of the case when the graph of $u_{0}$ is a grim reaper plane.
@@
It turns out that by choosing the blowup sequence in a particular way, these limit flows are self similar. For example, Huisken [undefv] showed that type I singularity formation can be modeled on self shrinking solutions. On the other hand, by following Hamilton’s blowup procedure in [undefp], Huisken and Sinestrari [undefx], [undefw] showed that type II singularity formation of mean convex mean curvature flow can be modeled on translating solutions.
@@
In this paper, we focus on the dynamical stability of various graphical translators. Stability of solitons is a widely studied subject, and tell us how ‘rigid’ the soliton is under perturbations. Stability is usually studied in two different contexts, linear stability and dynamical stability. In the setting of translators, Ilmanen [undefy] showed that translators are minimal hypersurfaces in Euclidean space under the conformal metric
@@
In [undefah], Xin showed that all graphical translators are area minimizing, hence linearly stable.
@@
We first look at the one dimensional mean curvature flow, also called curve shortening flow. Mullins [undefac] introduced a graphical translator in $\mathbf{R}^{2}$, which is now called the grim reaper. It turns out that this is the unique one dimensional translator up to rigid motion (see [undefo]). Wang and Wo [undefad] proved a dynamical stability result for grim reaper.
We generalize the result of Wang and Wo [undefad] by weakening the assumptions on the perturbation.
@@
Wang and Wo [undefad] assumed that $\|u_{0}-\overline{u}\|_{C^{2+\alpha}(-\frac{\pi}{2},\frac{\pi}{2})}<\infty$, and $M_{0}$ has finitely many inflection points. On the other hand, in theorem 1.3, we do not impose any form of convexity assumption. Also, one can easily see that finite curvature condition follows from
@@
Thus our assumptions are indeed weaker than the $C^{2+\alpha}$ closeness condition in [undefad].
@@
Theorem 1.3 implies that one cannot obtain an analogous result to theorem 1.1 in [undefj] for grim reapers. Even if $u_{0}\in C^{\infty}(-\frac{\pi}{2},\frac{\pi}{2})$ is identical to $\overline{u}$ for $x\neq(-\epsilon,\epsilon)$ for any small $\epsilon>0$, if $u_{0}>\overline{u}$ in $(-\epsilon,\epsilon)$, then $c_{0}>0$ in theorem 1.3, hence we do not have convergence to $\overline{M}+te_{n+1}=\text{Graph}(\overline{u}+t)$, but instead to its vertical translate.
We now consider surfaces in $\mathbf{R}^{3}$. Altschuler and Wu [undef] showed the existence of bowl soliton in $\mathbf{R}^{3}$, which is a rotationally symmetric, graphical translator. Hoffman et.al [undefu] constructed a one parameter family of strictly convex, graphical translators defined over a slab. In the same paper, by building on previous work of Spruck and Xiao [undefab], the authors also showed that in $\mathbf{R}^{3}$, the only graphical translators are the (tilted) grim reaper plane, 2d bowl soliton, and the one parameter family of translators they constructed, which are called $\Delta$-wings.
@@
We now consider higher dimensional translators. Clutterbuck et.al [undefj] proved the existence, uniqueness, and dynamical stability result for the bowl soliton in all dimensions. Later, Haslhofer [undefs] proved a uniqueness result of the bowl soliton among uniformly two convex, non-collapsed translators.
@@
Comparing with the dynamical stability result of Clutterbuck et.al [undefj], our theorem considers all possible non-collapsed translators. Moreover, we still obtain a stability result without a strong assumption on the asymptotic behavior of the initial graph at farfield. On the other hand, we require mean convexity, which can be viewed as a $C^{2}$ closeness condition.
@@
We explain the main strategy to proving dynamical stability of various graphical translators. Classical regularity theory for graphical solutions to mean curvature flow established in [undefn], [undefm], and theorem 1.1 implies the existence of longtime graphical solutions to mean curvature flow starting from the perturbed hypersurface. Let us denote the solution by $\{M_{t}\}_{t\in[0,\infty)}$. We then look at the forward limit of the flow. For any $t_{j}\to\infty$, we define
@@
The first task is to show that the limit flow is also a graphical translating solution with the same velocity $e_{n+1}$. For theorem 1.3, and theorem 1.6, this will be achieved by using the Harnack inequality of Hamilton [undefq]. For theorem 1.9, this will be done by showing that the limit flow we obtained is an asymptotically cylindrical flow. Then by using the recent classification result of [undefc], together with the fact that the limit flow is eternal, bounded between two translating solutions, we can conclude that the limit flow is a non-collapsed translator of the same velocity as the two barrier solutions. We then use existing classification theorems of graphical translators to conclude that the limit flow is actually equal to the translating solutions generated by the initial translator up to space translation.
@@
  • Uniqueness of one dimensional translators (see [undefo]).
  • Uniqueness of bowl solitons (see [undefs]).
  • Classification of graphical translators in $\mathbf{R}^{3}$ (see [undefu]).
@@ 2 Notations and preliminaries
We follow the notations in [undefn].
@@
Next, we recall Hamilton’s differential Harnack inequality for mean curvature flow [undefq].
@@ 3 Existence of graphical solution
In this section, we prove the existence of longtime graphical solution to mean curvature flow over a slab $\Omega^{n}_{b}=\mathbf{R}^{n-1}\times(-b,b)$ with continuous initial data. Following the ideas in [undefj], we construct a sequence of entire graph solutions to mean curvature flow with approximate initial data. We show that one can take a $C^{\infty}_{loc}(\Omega^{n}_{b})$ subsequential limit, which is the desired graphical solution over a slab. Due to standard quasilinear PDE theory, this can be done by obtaining an apriori $C^{1}$ estimate of graphical solutions over $\Omega^{n}_{b}$. $C^{0}$ estimate is obtained by using ancient pancakes of Bourni et.al [undefe] as barriers. Then, one can apply the gradient estimate of Colding - Minicozzi [undefk] to obtain the desired $C^{1}$ estimate.
@@
Once we have a $C^{0}$ bound, we can apply the gradient estimate in [undefk] to obtain a $C^{1}$ estimate. Let us first recall the gradient estimate in [undefk].
Lemma 3.3 (Theorem 1 in [undefk]).
@@
Once we have an interior gradient estimate, we can apply classical interior curvature estimates of Ecker and Huisken in [undefn] to obtain higher order derivative estimates to graphical solution $u$ by restricting it to small standard parabolic cylinders as we did in the proof of lemma 3.4. This implies the following result.
@@
Then, by classical results of Ecker and Huisken in [undefn], there exist sequence of functions $u^{j}\in C^{\infty}(\mathbf{R}^{n}\times[0,\infty))$ which solves the initial value problem
@@ 4 Dynamical stability of grim reaper
We generalize theorem 1.4 in [undefad] by weakening the initial $C^{2+\alpha}$ closeness (condition 1.10 in [undefad]), and dropping the finitely many inflection points condition (condition 1.11 in [undefad]). We also mention that our approach is different from that in [undefad], where the authors used the monotonicity of
@@
By assumption (89), (90), and the fact that $M_{0}$ is graphical solution over $(-\frac{\pi}{2},\frac{\pi}{2})$, we see that $M_{0}$ is uniformly $C^{1}$ close to vertical lines $\{x=\frac{\pi}{2}\}(\{x=-\frac{\pi}{2}))$ as $x\to\frac{\pi}{2}(x\to-\frac{\pi}{2})$. In other words, for any $\epsilon>0$, we can find $\delta_{0}>0$ so that for any $\frac{\pi}{2}-\delta_{0}<|x|<\frac{\pi}{2}$, we can write $M_{0}$ as a $C^{1}$ graph over vertical line with height and gradient bounded by $\epsilon>0$. Then by pseudolocality theorem (see theorem 1.5 in [undefz]), the same property (uniform $C^{1}$ closeness to vertical line) hold for all $M_{t}$, $t\in[0,T]$. Then by interior regularity theorem in [undefm] with assumption (90), we indeed have uniform bound on curvature up to $t=T$. Once we have bounded curvature and finite initial total curvature, we can apply lemma 6.1 in [undefai] to conclude that
@@
By assumption (89) and avoidance principle [undefaf], if we write $M^{j}_{t}=\text{Graph}(u^{j}(\cdot,t))$, then
@@
The limit flow is a complete, eternal, graphical solution to mean curvature with total curvature bounded by $I_{0}$, and lies in between two translating grim reapers. We claim that $M^{\infty}_{t}$ is convex. We follow the argument in proposition 5.4 in [undefai]. We claim that whenever $\kappa_{\infty}(p,t)=0$ at some $p\in M^{\infty}_{t}$, then $\partial_{s}\kappa_{\infty}(p,t)=0$ as well. Suppose this is not the case and there exists $p\in M^{\infty}_{t_{0}}$ so that
@@
Following the computations in lemma 6.1 in [undefai], for all $j\geq j_{0}$, $|t-t_{0}|<\delta$, $0<\epsilon<1$, and $N$ sufficiently large (but independently of the parameters $j,t,\epsilon$), we have
@@
This means that all inflection points of $M^{\infty}_{t}$ are degenerate. By using zero set result in [undefa] (see also proof of theorem 6.1 in [undefr]), this implies that $M^{\infty}_{t}$ is convex. It is actually strictly convex, or else by strong maximum principle, the limit flow has to be a straight line which is impossible since it lies between two translating grim reapers.
So far, we have seen that the limit flow has to be strictly convex. Also, by once again using finite total curvature and the fact that it lies between two grim reapers, we can obtain uniformly $C^{1}$ closeness of the limit flow to vertical lines $\{x=\frac{\pi}{2}\}(\{x=-\frac{\pi}{2}\})$ as $|x|\to\frac{\pi}{2}$. Then local regularity theorem [undefae] implies that the limit flow has local in time uniform bound on the curvature. Thus we can use Hamilton’s Harnack inequality (29) on the limit flow. As in proposition 6.1 of [undefg], we define the vector field
@@
By following [undefad], we now claim that

### v2

@@ 1 Introduction
Ecker and Huisken [undefo], [undefn] showed by maximum principle arguments that when the initial hypersurface $M_{0}$ is an entire graph of a locally Lipchitz function, then there exists a longtime smooth solution to mean curvature flow starting from $M_{0}$ which its time slices are also entire graphs. Furthermore, they showed that if $M_{0}$ is ‘straight at infinity’, then after suitable rescaling, the hypersurfaces converge to a self expander, a special type of hypersurfaces which move by expansion under mean curvature flow. Later, Clutterbuck [undefi] (see also [undefj]) relaxed local Lipchitz continuity to just continuity. Uniqueness of entire graph solutions starting from locally Lipchitz, proper graph was proved by Daskalopolous and Sáez [undefm].
In the case when $\Omega\subset\mathbf{R}^{n}$ is a general domain, Sáez and Schnürer [undefab] showed the existence of solutions to mean curvature flow when the initial data is a graph of a locally Lipchitz, proper function over $\Omega$. They further showed that such solutions remain graphical over evolving domain.
Our first result is the longtime existence of a graphical solution to mean curvature flow starting from a graph of a continuous function defined over $\Omega^{n}_{b}=\mathbf{R}^{n-1}\times(-b,b)$ for some $b>0$, analogous to results in [undefo], [undefi].
@@
Compared to results in [undefab], although we are only working with $\mathbf{R}^{n-1}\times(-b,b)$ instead of general domains, our result relaxes the local Lipchitz continuity to just continuity. We are also assuming a slightly weaker condition on $u_{0}$ than it being a proper function in view of the case when the graph of $u_{0}$ is a grim reaper plane.
@@
It turns out that by choosing the blowup sequence in a particular way, these limit flows are self similar. For example, Huisken [undefw] showed that type I singularity formation can be modeled on self shrinking solutions. On the other hand, by following Hamilton’s blowup procedure in [undefq], Huisken and Sinestrari [undefy], [undefx] showed that type II singularity formation of mean convex mean curvature flow can be modeled on translating solutions.
@@
In this paper, we focus on the dynamical stability of various graphical translators. Stability of solitons is a widely studied subject, and tell us how ‘rigid’ the soliton is under perturbations. Stability is usually studied in two different contexts, linear stability and dynamical stability. In the setting of translators, Ilmanen [undefz] showed that translators are minimal hypersurfaces in Euclidean space under the conformal metric
@@
In [undefaj], Xin showed that all graphical translators are area minimizing, hence linearly stable.
@@
We first look at the one dimensional mean curvature flow, also called curve shortening flow. Mullins [undefae] introduced a graphical translator in $\mathbf{R}^{2}$, which is now called the grim reaper. It turns out that this is the unique one dimensional translator up to rigid motion (see [undefp]). Wang and Wo [undefaf] proved a dynamical stability result for grim reaper.
We generalize the result of Wang and Wo [undefaf] by weakening the assumptions on the perturbation.
@@
Wang and Wo [undefaf] assumed that $\|u_{0}-\overline{u}\|_{C^{2+\alpha}(-\frac{\pi}{2},\frac{\pi}{2})}<\infty$, and $M_{0}$ has finitely many inflection points. On the other hand, in theorem 1.3, we do not impose any form of convexity assumption. Also, one can easily see that finite curvature condition follows from
@@
Thus our assumptions are indeed weaker than the $C^{2+\alpha}$ closeness condition in [undefaf].
@@
Theorem 1.3 implies that one cannot obtain an analogous result to theorem 1.1 in [undefj] for grim reapers. Even if $u_{0}\in C^{\infty}(-\frac{\pi}{2},\frac{\pi}{2})$ is identical to $\overline{u}$ for $x\notin(-\epsilon,\epsilon)$ for any small $\epsilon>0$, if $u_{0}>\overline{u}$ in $(-\epsilon,\epsilon)$, then $c_{0}>0$ in theorem 1.3, hence we do not have convergence to $\overline{M}+te_{n+1}=\text{Graph}(\overline{u}+t)$, but instead to its vertical translate.
We now consider surfaces in $\mathbf{R}^{3}$. Altschuler and Wu [undef] showed the existence of bowl soliton in $\mathbf{R}^{3}$, which is a rotationally symmetric, graphical translator. Hoffman et.al [undefv] constructed a one parameter family of strictly convex, graphical translators defined over a slab. In the same paper, by building on previous work of Spruck and Xiao [undefad], the authors also showed that in $\mathbf{R}^{3}$, the only graphical translators are the (tilted) grim reaper plane, 2d bowl soliton, and the one parameter family of translators they constructed, which are called $\Delta$-wings.
@@
We now consider higher dimensional translators. Clutterbuck et.al [undefj] proved the existence, uniqueness, and dynamical stability result for the bowl soliton in all dimensions. Later, Haslhofer [undeft] proved a uniqueness result of the bowl soliton among uniformly two convex, non-collapsed translators.
@@
One can see from the proof of theorem 1.9 that the mean convexity condition in theorem 1.9 is only used to exclude potential high multiplicity tangent flows. In particular, one can impose different assumption in place of (ii), and still get the same result. In particular, we have the following corollary of the proof of theorem 1.9.
Corollary 1.12.
Following the notations in theorem 1.9, suppose
$(i)\ \|u_{0}-\overline{u}\|_{C^{0}(\mathbf{R}^{n})}\leq C_{0}<\infty,\ \ \ (ii)\ \lambda(M_{0})<2\lambda(\overline{M}),$ (16)
where $\lambda$ denotes the Colding-Minicozzi entropy (see definition 2.6). Then the same conclusion as theorem 1.9 holds.
Remark 1.13.
Comparing with the dynamical stability result of Clutterbuck et.al [undefj], our theorem considers all possible non-collapsed translators. Moreover, we still obtain a stability result without a strong assumption on the asymptotic behavior of the initial graph at farfield. On the other hand, we require mean convexity, which can be viewed as a $C^{2}$ closeness condition. Indeed, by the translator equation (3) written using $\overline{u}$, we have
$Q(D\overline{u},D^{2}\overline{u})=(\delta_{ij}-\frac{D_{i}\overline{u}D_{j}\overline{u}}{1+|D\overline{u}|^{2}})D^{2}_{ij}\overline{u}=1.$ (17)
Since $Q$ is smooth with respect to its arguments, we see that even after perturbing $D\overline{u},D^{2}\overline{u}$ by a small amount, $Q$ will still be positive, meaning the perturbed surface is still mean convex.
@@
We explain the main strategy to proving dynamical stability of various graphical translators. Classical regularity theory for graphical solutions to mean curvature flow established in [undefo], [undefn], and theorem 1.1 implies the existence of longtime graphical solutions to mean curvature flow starting from the perturbed hypersurface. Let us denote the solution by $\{M_{t}\}_{t\in[0,\infty)}$. We then look at the forward limit of the flow. For any $t_{j}\to\infty$, we define
@@
The first task is to show that the limit flow is also a graphical translating solution with the same velocity $e_{n+1}$. For theorem 1.3, and theorem 1.6, this will be achieved by using the Harnack inequality of Hamilton [undefr]. For theorem 1.9, this will be done by showing that the limit flow we obtained is an asymptotically cylindrical flow. Then by using the recent classification result of [undefc], together with the fact that the limit flow is eternal, bounded between two translating solutions, we can conclude that the limit flow is a non-collapsed translator of the same velocity as the two barrier solutions. We then use existing classification theorems of graphical translators to conclude that the limit flow is actually equal to the translating solutions generated by the initial translator up to space translation.
@@
  • Uniqueness of one dimensional translators (see [undefp]).
  • Uniqueness of bowl solitons (see [undeft]).
  • Classification of graphical translators in $\mathbf{R}^{3}$ (see [undefv]).
@@ 2 Notations and preliminaries
We follow the notations in [undefo].
@@
Next, we recall Hamilton’s differential Harnack inequality for mean curvature flow [undefr].
@@
Next, we recall the definition of Colding-Minicozzi entropy, which was first introduced in [undefk]. We state the definition for Radon measures in $\mathbf{R}^{n+1}$. One can then understand the entropy of a hypersurface $M$ as the entropy of the Radon measure $\mathcal{H}^{n}|_{M}$, where $\mathcal{H}^{n}$ is the $n$-dimensional Hausdorff measure.
Definition 2.6.
Let $\mu$ be a Radon measure in $\mathbf{R}^{n+1}$ which satisfies
$\sup_{r>0}\frac{\mu(B(0,r))}{r^{n}}<\infty.$ (32)
Then for each $x_{0}\in\mathbf{R}^{n+1}$, $t>0$, define
$F_{x_{0},t}(\mu)=\int_{\mathbf{R}^{n+1}}\frac{1}{(4\pi t)^{n/2}}\exp{(-\frac{|x-x_{0}|^{2}}{4t})}d\mu.$ (33)
Then, the entropy of $\mu$ is defined as
$\lambda(\mu)=\sup_{x_{0}\in\mathbf{R}^{n+1},t>0}F_{x_{0},t}(\mu).$ (34)
Two key facts about entropy is that it is invariant under translation / dilation, and it is non-increasing under mean curvature flow.
@@ 3 Existence of graphical solution
In this section, we prove the existence of longtime graphical solution to mean curvature flow over a slab $\Omega^{n}_{b}=\mathbf{R}^{n-1}\times(-b,b)$ with continuous initial data. Following the ideas in [undefj], we construct a sequence of entire graph solutions to mean curvature flow with approximate initial data. We show that one can take a $C^{\infty}_{loc}(\Omega^{n}_{b})$ subsequential limit, which is the desired graphical solution over a slab. Due to standard quasilinear PDE theory, this can be done by obtaining an apriori $C^{1}$ estimate of graphical solutions over $\Omega^{n}_{b}$. $C^{0}$ estimate is obtained by using ancient pancakes of Bourni et.al [undefe] as barriers. Then, one can apply the gradient estimate of Colding - Minicozzi [undefl] to obtain the desired $C^{1}$ estimate.
@@
Once we have a $C^{0}$ bound, we can apply the gradient estimate in [undefl] to obtain a $C^{1}$ estimate. Let us first recall the gradient estimate in [undefl].
Lemma 3.3 (Theorem 1 in [undefl]).
@@
Once we have an interior gradient estimate, we can apply classical interior curvature estimates of Ecker and Huisken in [undefo] to obtain higher order derivative estimates to graphical solution $u$ by restricting it to small standard parabolic cylinders as we did in the proof of lemma 3.4. This implies the following result.
@@
Then, by classical results of Ecker and Huisken in [undefo], there exist sequence of functions $u^{j}\in C^{\infty}(\mathbf{R}^{n}\times[0,\infty))$ which solves the initial value problem
@@ 4 Dynamical stability of grim reaper
We generalize theorem 1.4 in [undefaf] by weakening the initial $C^{2+\alpha}$ closeness (condition 1.10 in [undefaf]), and dropping the finitely many inflection points condition (condition 1.11 in [undefaf]). We also mention that our approach is different from that in [undefaf], where the authors used the monotonicity of
@@
By assumption (94), (95), and the fact that $M_{0}$ is graphical solution over $(-\frac{\pi}{2},\frac{\pi}{2})$, we see that $M_{0}$ is uniformly $C^{1}$ close to vertical lines $\{x=\frac{\pi}{2}\}(\{x=-\frac{\pi}{2}))$ as $x\to\frac{\pi}{2}(x\to-\frac{\pi}{2})$. In other words, for any $\epsilon>0$, we can find $\delta_{0}>0$ so that for any $\frac{\pi}{2}-\delta_{0}<|x|<\frac{\pi}{2}$, we can write $M_{0}$ as a $C^{1}$ graph over vertical line with height and gradient bounded by $\epsilon>0$. Then by pseudolocality theorem (see theorem 1.5 in [undefaa]), the same property (uniform $C^{1}$ closeness to vertical line) hold for all $M_{t}$, $t\in[0,T]$. Then by interior regularity theorem in [undefn] with assumption (95), we indeed have uniform bound on curvature up to $t=T$. Once we have bounded curvature and finite initial total curvature, we can apply lemma 6.1 in [undefak] to conclude that
@@
By assumption (94) and avoidance principle [undefah], if we write $M^{j}_{t}=\text{Graph}(u^{j}(\cdot,t))$, then
@@
The limit flow is a complete, eternal, graphical solution to mean curvature with total curvature bounded by $I_{0}$, and lies in between two translating grim reapers. We claim that $M^{\infty}_{t}$ is convex. We follow the argument in proposition 5.4 in [undefak]. We claim that whenever $\kappa_{\infty}(p,t)=0$ at some $p\in M^{\infty}_{t}$, then $\partial_{s}\kappa_{\infty}(p,t)=0$ as well. Suppose this is not the case and there exists $p\in M^{\infty}_{t_{0}}$ so that
@@
Following the computations in lemma 6.1 in [undefak], for all $j\geq j_{0}$, $|t-t_{0}|<\delta$, $0<\epsilon<1$, and $N$ sufficiently large (but independently of the parameters $j,t,\epsilon$), we have
@@
This means that all inflection points of $M^{\infty}_{t}$ are degenerate. By using zero set result in [undefa] (see also proof of theorem 6.1 in [undefs]), this implies that $M^{\infty}_{t}$ is convex. It is actually strictly convex, or else by strong maximum principle, the limit flow has to be a straight line which is impossible since it lies between two translating grim reapers.
So far, we have seen that the limit flow has to be strictly convex. Also, by once again using finite total curvature and the fact that it lies between two grim reapers, we can obtain uniformly $C^{1}$ closeness of the limit flow to vertical lines $\{x=\frac{\pi}{2}\}(\{x=-\frac{\pi}{2}\})$ as $|x|\to\frac{\pi}{2}$. Then local regularity theorem [undefag] implies that the limit flow has local in time uniform bound on the curvature. Thus we can use Hamilton’s Harnack inequality (31) on the limit flow. As in proposition 6.1 of [undefg], we define the vector field
@@
By following [undefaf], we now claim that